OpenStudy (anonymous):

use mental math to solve each equation... 1. 2^x=1/8 2. log (lowered 4) 2=x 3. 10 ^6^x =1 4. 5^2^x = 125

OpenStudy (anonymous):

i'm so lost

OpenStudy (luigi0210):

for 1 you could rewrite it as:2^x=2^-3

OpenStudy (luigi0210):

:/

OpenStudy (anonymous):

would that solve it though?

OpenStudy (luigi0210):

yes

OpenStudy (anonymous):

okay I'll check once I finish the homework!

OpenStudy (anonymous):

what would log5/3log4 be? I don' t have a calculator that has log on it?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

what problem are you doing..?

OpenStudy (anonymous):

4^3^n=16

OpenStudy (anonymous):

i have to solve it so i have to break it down and then again through power property and divide log5/3log4 but on the online calculator I was using it gave me .14 and thats not the answer :/

OpenStudy (luigi0210):

\[4^{3^n}=4^2\] \[3^n=2\]

OpenStudy (luigi0210):

Is that how you started?

OpenStudy (anonymous):

no,